# A non-uniform metal distribution in the ram-pressure stripped M86 group of galaxies

### v1

@@ 4.2.2 SNIa explosions: Discrepancy between SNIa models based on their explosion symmetry
Our results in Table 4.2 indicate a discrepancy in the SNIa contributions between 3D pure deflagration and delayed-detonation models if the explosion is asymmetric (i.e., a low number of ignition spots). We see that, based on our 5 best-fitting results, the required SNIa contribution is between 59–88% for Fink models, and 15–25% for Seitenzahl models. Similarly, if we limit the IMF Slope to be only Salpater IMF (i.e., ), the same discrepancy between burning mechanisms is still prominent for 5 best-fitting results obtained from a Salpater IMF.
@@
Although our results suggest asymmetric SNIa explosions, we caution the reader that the absence of key SNIa products, such as Ar, Ca, and Ni, could introduce biases and lead to poor constraints on the SNIa contribution. An abundance pattern of M86 with the addition of SNIa products such as Ar, Ca, and Ni may result in different best-fitting results, which could indicate normal SNIa explosions with a higher number of ignition spots, similar to those observed in closed-box galaxy clusters. The robust nature of SNe explosions in M86 can be revealed with future observations from the new-generation XRISM and future-generation NewAthena observatories, which are capable of accurately measuring SNIa products.
@@ 4.2.3 SNe and AGB yields in the M86 core
We find that the abundance pattern in the M86 core is well explained by non-rotating, low-metallicity SNcc progenitors, and geometrically asymmetric deflagration and delayed-detonation SNIa models. We see that delayed-detonation Seitenzahl models are more consistent with varying model parameters compared to deflagration Fink models. When using a Nomoto SNcc model and a Salpeter IMF, the delayed-detonation scenario—preferred by the supernova community—yields an SNIa contribution of 15–21%, based on Seitenzahl models. Allowing the IMF slope to vary, we find that the SNIa contribution remains consistent, ranging between 15–25%.
@@
Additionally, after fitting the SNIa and SNcc models to our abundance pattern of SNe products, we incorporated asymptotic giant branch (AGB) star models to fit our N measurement as well. AGB stars produce light elements like N at the end of their life and enrich their surroundings through stellar winds. To account for the combined contributions of AGB stars, SNcc, and SNIa, we used AGB models from Nomoto et al. (2013), adapted from Karakas (2010), and refit our data with the inclusion of the N/Fe ratio.
@@
We assumed the initial metallically is the same for SNcc progenitors and AGB stars. In our best-fitting combination, we used Nomoto SNcc model, Salpeter IMF; and Solar initial metallicity () for both SNcc progenitors and AGB stars. For the SNIa component, we used the Seitenzahl model with a single ignition spot (); the result is presented in the left panel of Fig. 12, where the colored bars represent the contributions of AGB or SNe production based on the given combination of models, while the black data points correspond to our measured M86 abundance ratios. In the same figure, we also present our result obtained from fit, in which Fink and Seitenzahl results agree.
4.3 The dissimilar metal distribution in the hot gas of M86
4.3.1 Mg/Fe ratio
The Mg/Fe ratios in Plume and Tail 1 are in agreement with that of M86 group outskirts and Virgo ICM. Therefore the Plume and Tail structures might have formed from this aforementioned mixed low-entropy group gas after the collision with NGC 4438 and the onset of ram-pressure stripping. Since these regions extend well beyond the stellar population of the galaxy and experience an accretion cut-off from their surroundings due to ram-pressure stripping (Gunn & Gott, 1972), the chemical composition in emission structures might have remained unchanged since their formation. Similarly, the chemical composition of the extended group gas (Regions 1 and 2), which is mixed efficiently with the Virgo ICM, is expected to preserve its chemical composition compared to that of the galaxy core. In contrast, in the M86 galaxy core, SNIa events continue to enrich the medium. Therefore, SNIa products such as Fe are expected to increase in the galaxy core, leading to a decreased Mg/Fe ratio.
Additionally, ram-pressure stripping and galaxy-galaxy interactions enhance the star formation, which increases the number of massive stars and SNcc explosions. In Tail 2 (Region 8), the far end of the M86 tail in our observation, we observe the highest Mg/Fe ratio in M86. It is plausible that the dynamical disruption in the M86 tail enhanced the Mg enrichment.
4.3.2 Si/Fe and S/Fe ratios
4.3.3 Origins of the emission structures
The chemical compositions of the M86 plume and tail show significant deviations from that of the galaxy core, and the Mg/Fe ratio in the plume is remarkably similar to that in the M86 group outskirts and the Virgo ICM. If these structures had formed primarily through the stripping of dense core gas, their chemical composition would be expected to be comparable. Similarly, the emission structures have been found to have low entropy (Finoguenov et al., 2004). We also showed that the plume and the tail, despite their considerably large extraction regions, are reasonably isothermal, with a Gaussian temperature distribution characterized by a width of of the mean temperature . If these emission structures had recently formed from the stripping of the dense galaxy core, their temperature structure would be expected to show variations due to ongoing mixing of the core gas. Their narrow temperature distributions indicate that these structures have had sufficient time to reach thermal equilibrium.
Our results suggest that the emission structures possibly originate from the initial low-entropy group gas of M86 which has efficiently mixed with the surrounding Virgo ICM gas, rather than the gas within the M86 galaxy core. This supports the argument that the galaxy-galaxy interaction is the primary cause of the emission structures, rather than ram-pressure stripping of the galaxy core.
@@ 4.3.3 Origins of the emission structures
Additionally, the absolute abundances and abundance ratios within the north-eastern arm are quite peculiar. We observe that this structure has the lowest absolute abundances for all elements. It shows a Mg/Fe ratio comparable to that in the core but a Si/Fe ratio more than higher than in the core. The S/Fe ratio in the north-eastern arm agrees with the S/Fe ratio everywhere except in the core.
@@ 5 Conclusions
Our analysis of supernova explosion models, based on the abundance pattern in the core of the M86 galaxy, suggests that the hot atmosphere may have been enriched by an early generation of SNcc explosions from a low-metallicity environment, with non-rotating SNcc progenitors, despite the galaxy-galaxy interaction and ram-pressure stripping. This result is consistent with those observed in closed-box galaxy clusters.
The emission structures of plume and tail have significantly different Mg/Fe ratio compared to the galaxy core. The plume structure has a Mg/Fe ratio 3.3 higher than the core, while our two tail regions show 2.5 and 4.3 higher Mg/Fe than the core. On the other hand, the Mg/Fe in these emission structures are remarkably comparable with that of the outskirts of the M86 group and also with the Virgo ICM. Therefore, it is possible to conclude that the plume and tail structures did not originate from the hot gas in the galaxy core through stripping, but rather from the group gas, which has well-mixed with the Virgo ICM. We note that the second tail region shows higher Mg/Fe ratios compared to the Virgo ICM and M86 group halo, possibly due to the dynamical interactions induced enhanced star-formation and SNcc events.
@@
We conclude that based on their chemical composition, the emission structures may originate from the disruption of the low-entropy group gas during the collision of NGC 4434, rather than the ram-pressure stripping of the dense galaxy core gas.
@@ Acknowledgements
SK would like to thank Boğaziçi University BAP, project code 19951, for their support. NW is supported by the GACR EXPRO grant No. 21-13491X. ENE would like to thank Boğaziçi University BAP, project code 13760, for their support.

### v2

@@ 4.3.2 SNIa explosions: Discrepancy between SNIa models based on their explosion symmetry
Our results in Table 4.3 indicate a discrepancy in the SNIa contributions between 3D pure deflagration and delayed-detonation models if the explosion is asymmetric (i.e., a low number of ignition spots). We see that, based on our 5 best-fitting results, the required SNIa contribution is between 59–88% for Fink models, and 15–25% for Seitenzahl models. Similarly, if we limit the IMF Slope to be only Salpeter IMF (i.e., ), the same discrepancy between burning mechanisms is still prominent for 5 best-fitting results obtained from a Salpeter IMF.
@@
Based on our SNIa contribution results over a wide range of progenitor parameters, we observe the following: (i) When SNcc progenitor parameters are allowed to vary freely, based on the best fits, the 3D delayed-detonation models –which are preferred by the supernova community– result in consistently low SNIa and high SNcc contributions, for both Nomoto and Limongi. (ii) When SNcc progenitor parameters are fixed and only the SNIa model parameters are varied, the delayed-detonation and deflagration models converge at an SNcc contribution of . However, the fit quality is poor in this case for the delayed-detonation model, and acceptable fits are only achieved when the SNcc contribution exceeds .
@@ 4.3.3 SNe and AGB yields in the M86 core
After fitting the SNIa and SNcc models to our abundance pattern of SNe products, we incorporated asymptotic giant branch (AGB) star models to fit our N measurement as well. AGB stars produce light elements like N at the end of their life and enrich their surroundings through stellar winds. To account for the combined contributions of AGB stars, SNcc, and SNIa, we used AGB models from Nomoto et al. (2013), adapted from Karakas (2010), and refit our data with the inclusion of the N/Fe ratio.
@@
We assumed the initial metallically is the same for SNcc progenitors and AGB stars. In our best-fitting combination, we used Nomoto SNcc model, Salpeter IMF; and Solar initial metallicity () for both SNcc progenitors and AGB stars. For the SNIa component, we used the Seitenzahl model with a single ignition spot (); the result is presented in Fig. 12, where the colored bars represent the contributions of AGB or SNe production based on the given combination of models, while the black data points correspond to our measured M86 abundance ratios. We caution the reader that our approximations assume that the SNe models with fixed configurations from our best-fits dominate the overall enrichment, while in reality, it may involve different generations of stars and configurations. Nonetheless, our best-fitting results suggest a high SNcc enrichment in M86 galaxy core with a SNcc contribution of 75–85% based on 3D delayed-detonation SNIa models.
4.4 Origins of the emission structures: ram-pressure stripping or galaxy-galaxy interaction?
In M86, we observe an (i) X-ray emitting tail with a true length of >380 kpc, which is almost twice the projected length, according to the orbital estimations of Randall et al. (2008), (ii) a plume at the north without any stellar counterparts which has the highest absolute abundances in our analysis, and (iii) a northeastern structure with the lowest absolute abundances. Formation histories of these emission structures in M86 are still in debate. With our elemental abundance results, we aim to constrain the possible scenarios in this section. The ram-pressure stripping occurs when the ram-pressure surpasses the thermal pressure. M86 group enters into Virgo Cluster with velocity and the Virgo ICM has electron density at 350 kpc from the center of Virgo Cluster of (Urban et al., 2011). Therefore we estimate the ram-pressure applied to M86 to be . Similarly, assuming the electron density of M86 core (Randall et al., 2008), the thermal pressure is found to be . Since , the hot halo of M86 experiences strong ram-pressure stripping as it moves supersonically in Virgo ICM.
Furthermore, the neighboring galaxy NGC 4438, located east of M86 with a projected separation of 120 kpc, is a large spiral galaxy that also exhibits strong morphological disturbances, particularly in its outer stellar disk, where extraplanar tidal arms indicate a past gravitational interaction. Both NGC 4438 and M86 are blueshifted with respect to M87, infalling toward the Virgo cluster from the far side, with clustercentric line-of-sight velocities of approximately and , respectively. Kenney et al. (2008) reported a projected 120 kpc H bridge between NGC 4438 and M86, and using this distance they estimate an average plane-of-sky velocity of for NGC 4438, and a timescale of Myr since the closest encounter with M86. In the rest frame of M86, they propose a collision scenario in which NGC 4438 enters the M86 system from the front with a relative line-of-sight velocity of and a transverse velocity of toward the east.
Based on the ram-pressure stripping and galaxy-galaxy collision histories of M86, combined with our results on the thermal and chemical structures in M86, we constrain the formation and evolution scenarios of the three emission structures in the following sections.
4.4.1 Tail
The X-ray tail of M86 extends to the northwest—exactly opposite the direction implied by any collision with NGC 4438. Furthermore, from the true length of the tail and velocity of M86, Randall et al. (2008) estimate a lower limit of the formation timescale of Myr. Since the encounter with NGC 4438 occurred only Myr ago, the tail must have been in place long before that collision. Thus, the tail is most naturally explained by ram-pressure stripping of M86’s hot halo as it moves through the Virgo ICM. Nevertheless, a subsequent interaction between the ram-pressure stripped tail of M86 and NGC 4438 may explain why the X-ray morphology of this tail differs from that of other elliptical galaxies in the Virgo and Fornax clusters undergoing ram-pressure stripping, which typically exhibit sharp surface brightness edges aligned with centrally trailing tails (Machacek et al., 2005, 2006). Furthermore, Finoguenov et al. (2004) showed that the tail is low-entropy. Similarly, we showed that the tail regions, despite their considerably large extraction areas, are reasonably isothermal, with a Gaussian temperature distribution characterized by a width of of the mean temperature . The narrow temperature distributions suggest that the tail has had sufficient time to reach thermal equilibrium.
The facts that the Mg/Fe and S/Fe ratios in the tail are significantly different from those in the core, along with the absence of temperature variations expected from the mixing of continuously stripped core gas, seem to contradict the simple stripping scenario. However, we remind the reader that based on the orbital calculations in Randall et al. (2008) in which M86 enters into Virgo cluster towards the observer and from northwestern direction, the lower limit on the actual length of the extended tail is more than twice the projected length. Therefore, tail sections that are spatially closer to the core—where mixing is expected—are very difficult to constrain in the projected image of M86, in which the plume is also another dominant emission source. Moreover, we note that it is also possible that, in addition to the galaxy core, the plume is being stripped and contributes to the formation of the visible tail. Thus, our Tail 1 and Tail 2 regions, which have large extraction areas, represent integrated stripped gas over a large physical volume and a long formation timescale, allowing them to reach thermal equilibrium.
Notably, the tail regions show increased Mg/Fe ratios compared to the plume and the galaxy core. Furthermore, we observe that the more distant tail region, Tail 2, shows enhanced Mg/Fe, Si/Fe and S/Fe ratios compared to the closer tail region, Tail 1. This result is in agreement with the ram-pressure stripping of gas and the consequent increase in star formation in the turbulent tail regions, accompanied by the increased SNcc explosions and elements.
4.4.2 Plume
In this study, we have shown that the plume in M86 has the highest absolute abundances for all elements in our measurements and the Mg/Fe ratio in the plume is in excellent agreement with that of the outer, less dense group gas and the Virgo ICM, while it significantly deviates from the Mg/Fe ratio in the core. Since M86 does not harbour an AGN, the possible mechanisms to create the plume structure are the ram-pressure stripping, the galaxy-galaxy interaction, the SNe explosions inside the galaxy core due to a starburst, and/or the combinations of these hot gas perturbations. A more exact computation of the impact of these mechanisms would require 3D simulations, which is out of scope of this paper. Instead, we perform simpler order of magnitude comparisons and together with our elemental abundance results we aim to constrain these scenarios. First, let us assume a scenario in which the gas is uplifted to a distance of 20 kpc from the galaxy core. The energy required for a such lifting is,
where is the initial average distance to the galaxy center of the uplifting gas. Note that this energy equation is oversimplified for an order of magnitude comparison and assumes all uplifting gas is initially concentrated as a shell at and the enclosed mass is constant. If we assume the half of the effective radius of M86 as the average initial distance, which is kpc, we find = using the total M86 enclosed mass at effective radius from Cappellari et al. (2013) and the hot gas mass of the plume from Randall et al. (2008). The other energy components of such displacement are the thermal heating energy due to slightly higher temperature of plume compared to the core, and the kinetic energy in the rest frame of M86 galaxy core based on the 20 kpc distance and the fact that there is no shock fronts related to the plume and the motion is subsonic. However, we find these components to be one order of magnitude less than , therefore we ignore these terms.
If we assume the galaxy-galaxy collision or ram-pressure stripping induced a starburst in the galaxy core, it is expected that SNcc events and stellar winds eject material into the surroundings. In order to test the pure SNe scenario of the plume origin, let us assume the most extreme case, in which all the ejected material is placed at north of M86 core. We adopt the mechanical energy injection rate from SNcc and stellar winds as given by Veilleux et al. (2005) to compare this feedback energy with the gas lifting requirement. Assuming a star formation rate of 1, a duration of 100 Myr, and a thermalization efficiency of 10%, we find that the total mechanical energy output of such a starburst is , which is not comparable with the uplifting energy. We see that the origin of the plume cannot be explained by such a starburst alone. Nevertheless, M86 orbits toward the southeast direction, and this gas flow can displace the significant amount of ejected material to the north, enriching the environment. Therefore, this ram-pressure and/or galaxy-galaxy interaction induced starburst can explain the high absolute abundances in the plume.
As for the ram-pressure stripping scenario of the plume, Randall et al. (2008) report that stripping such a single blob is not possible with a spherical gravitational potential of M86, but it is only possible with an aspherical potential. In our study we find that, similar to the tail, the plume is remarkably isothermal and low in entropy. If the X-ray bright plume had recently formed from the stripping of the dense galaxy core and continued to receive inflow from the core, its temperature structure would be expected to show variations due to ongoing mixing of the core gas. Moreover, the Mg/Fe ratios in the plume and core are significantly different. Therefore, we conclude that the plume has not originated recently from the stripping of the dense core. Nevertheless, if the plume structure formation started at the same time as the tail as M86 enters into Virgo cluster as Randall et al. (2008) suggested, it should be formed from the same CGM gas of M86 as the tail 200–300 Myr ago. Therefore, if the plume has originated from the CGM of that period, it is expected to preserve its chemical composition while the galaxy core has continuously been enriched by SNIa products. The Mg/Fe ratios in the Virgo ICM, outer M86 hot gas, and the plume are Solar. The ratio in the core, on the other hand, is Solar. Therefore, in principle, the ongoing Fe production in the galaxy core might change the "initial" Mg/Fe ratio of Solar to a Solar over time. However, this scenario is entirely based on a "Fe excess" in the galaxy core due to ongoing SNIa events. The facts that the plume and the core has comparable hot gas masses (Randall et al., 2008) and that we observe times more Fe absolute abundance in the Plume compared to the Core is in contrast with this scenario101010We note that our Core region is considerably smaller than the core region used by Randall et al. (2008) for their mass determination. However, our Region 3, which is a circular core region that surrounds and excludes the inner Core region, is roughly similar in size to theirs, and the Fe abundances in our Region 3 and Core region are consistent within 1 ( and respectively). Therefore, the interpretation that the core contains less Fe mass than the plume remains valid..
The galaxy–galaxy collision scenario provides a natural explanation for the increased Mg/Fe ratio observed in the plume. A strong perturbation from the close passage of NGC 4438, combined with the turbulent conditions already present due to ram-pressure stripping, could effectively displace and compress M86’s CGM. Kenney et al. (2008) estimate the kinetic energy of the gas stripped from NGC 4438 during its interaction with M86’s hot halo to be , a value comparable to the total thermal energy of the M86 CGM. This level of energy, combined with ram-pressure flows, can facilitate the removal of the outer halo gas with super-Solar Mg/Fe ratio. Therefore, the observed chemical composition of the plume can be understood as the displaced, Mg rich outer halo gas. The caveat of this scenario is also the high absolute abundances in the plume, which is not observed in the outer halo gas. However, high absolute abundances can be explained by the gas loss from NGC 4438 during its passage and an increased starburst. We can speculate that the interaction of M86 halo with NGC 4438 ISM can induce a starburst and a subsequent SNcc enrichment in both NGC 4438 and M86, which can enrich the M86 halo medium during the passage. If the galaxy-galaxy collision and ram-pressure induced turbulent environment caused the formation of the plume, the plume might act like a repository for these ejected metals. Our high SNcc contribution result in Sect. 4.3 supports this argument.
In conclusion, a pure ram-pressure stripping scenario is unable to explain the high absolute abundances in the plume. The combination of ram-pressure stripping and galaxy-galaxy collision with NGC 4438 can explain the high absolute abundances and elevated Mg/Fe ratios in the plume, under the assumptions of the plume has originated from the perturbation of the outer CGM of M86 and further enriched by the starbursts in M86 and NGC 4438 due to ram-pressure, which is supported by our high SNcc contribution results. We note that tailored simulations are necessary to completely understand the plume phenomena in M86.
4.4.3 The north-eastern arm
@@
The north-eastern arm spatially corresponds with the H emission structures and the M86 edge of the H bridge between NGC 4438. The absolute abundances and abundance ratios within the north-eastern arm are quite peculiar. We observe that this structure has the lowest absolute abundances for all elements. It shows a Mg/Fe ratio comparable to that in the core but a Si/Fe ratio more than higher than in the core. The S/Fe ratio in the north-eastern arm agrees with the S/Fe ratio everywhere except in the core. It is possible that this region, which is not isothermal like the plume and tail, is highly disturbed due to both galaxy-galaxy collision and ram-pressure stripping, and continuously mixed with the surrounding material.
@@ 5 Conclusions
Our analysis of supernova explosion models, based on the abundance pattern in the core of the M86 galaxy, suggests that the SNcc contribution to the hot gas enrichment is , higher than the closed-box galaxy clusters. Our SNcc contribution result suggests that the ram-pressure stripping and galaxy-galaxy interaction possibly increased the star formation rate and the consequent SNcc explosions.
The emission structures of plume and tail have significantly different Mg/Fe ratio compared to the galaxy core. The plume structure has a Mg/Fe ratio 3.3 higher than the core, while our two tail regions show 2.5 and 4.3 higher Mg/Fe than the core. On the other hand, the Mg/Fe in the plume is remarkably comparable with that of the outskirts of the M86 group and also with the Virgo ICM. Therefore, it is possible to conclude that the plume structure did not originate from the hot gas in the galaxy core through stripping, but rather from the group gas due to galaxy-galaxy interaction with NGC 4438. We note that the second tail region shows higher Mg/Fe ratios compared to the Virgo ICM and M86 group halo, possibly due to the dynamical interactions induced enhanced star-formation and SNcc events.
@@
We conclude that, based on their chemical composition, neither a pure galaxy–galaxy interaction nor pure ram-pressure stripping alone can explain the emission structures. While the tail possibly originates from ram-pressure stripping, the plume may result from the disruption of low-entropy outer hot gas during the collision of NGC 4438 combined with ram-pressure flows, rather than from the ram-pressure stripping of the dense galaxy core gas.
@@ Acknowledgements
The authors sincerely thank the anonymous referee for their thoughtful and constructive feedback, which improved the quality of this paper. SK would like to thank Boğaziçi University BAP, project code 19951, for their support. NW is supported by the GACR EXPRO grant No. 21-13491X. ENE would like to thank Boğaziçi University BAP, project code 13760, for their support.